KyledaGreat:

The total mechanical energy of an object with mass m at height h in a closed system can be written as ME=1/2mv^2+mgh. Solve for v, the velocity of the object, in terms of the given quantities.

KyledaGreat:

1 attachment
KyledaGreat:

@vocaloid

Vocaloid:

subtract mgh from both sides from there, you'll have (1/2)mv^2 on the right side divide both sides by (1/2)m and take the square root

KyledaGreat:

i'm sorry , i'm having difficulty in doing it

Vocaloid:

ME = (1/2)mv^2 + mgh subtracting mgh from both sides: ME - mgh = (1/2)mv^2 divide both sides by (1/2)m 2(ME - mgh)/m = v^2 take the square root of both sides v = sqrt[2(ME - mgh)/m]